the price elasticity of demand measures the responsiveness of the change in the

 22. The price elasticity of demand measures the responsiveness of the change in the:

 A) quantity demanded to a change in the price.

 B) price to a change in the quantity demanded.

 C) lope re enterprise

 D) slope of the demand curve to a change in the quantity demanded.


 23. The price of gasoline rises 5% and the quantity of gasoline purchased falls 1%.

 price elasticity of demand is equal to _______  and demand is described as _______ 

 A) 0.2; inelastic

 B) 5; inelastic

 C) 0.2; elastic


 24. For a normal demand curve, the price elasticity of demand will be:

 A) always positive.

 B) always greater than 1.

 C) usually equal to 1.

 D) always negative.


 25. A men's tie store sold an average of 30 ties per day at S5 per tie but sol

 ties per day at S3 per tie. The price elasticity of demand, by the mid

 A) greater than zero but less than 1.

 B) equal to 1.

 C) greater than 1 but less than 3.

 D) greater than 3.


 26. If the price of chocolate-covered peanuts decreases from $1.10 to SO...

 demanded increases from 190 bags to 210 bags, then the price elasticity,

 the midpoint method) is:

 A) 0.

 B) 0.5.

 C) 1.

 D) 2.
Answer #1

Q 22 (a)- price elasticity of demand measures the responsiveness of the change in quantity demanded to a change in price.

Q 23 (a) Elasticity is measured as the negative of ratio of percentage change in quantity to percentage change in price, So elasticity equals 1/5=0.2. If elasticity is less than 1 in absolute value we say that demand is inelastic.

Q 24 (d) by a normal demand curve we mean a downward sloping demand curve which means change in price and quantity move in opposite direction. So elasticity for a normal demand curve is always negative. But however note that the general convention is take absolute value as the elasticty.

Q 26 (b) In the midpoint formula of elasticity the idea is that in claculating the percentage change in quantity or prices we divide the difference between the new price or quantity and old price or quantity by the average of the new and old price or quantity (rather than than the old price or quantity which is generally used to calculate percentage).

%change in quantity= (190-210)/200=10%

%change in price= (1.1-0.90)/1=20%

thus elasticity =10/20=0.5

Note that your question 25 is not clear in the picture.Please re up;oad a new image of the question.